OpenStudy (anonymous):

Write an equation in slope-intercept form, point-slope, or standard form for the line with the given information. Explain why you chose the form you used. b. Slope 2, y-intercept -4

OpenStudy (anonymous):

slope intercept form is y = mx + b. m is the slope and b is the y-intercept. So just put 2 in for m and -4 in for b and there's your equation

OpenStudy (anonymous):

cool thanks can i have help with another one?

OpenStudy (anonymous):

sure

OpenStudy (anonymous):

c. Has an x-intercept of 6 and a y-intercept of 3

OpenStudy (anonymous):

the x intercept means that the line will cross x at 6 so you can put x = 6. Same with the y-intercept. It will cross y at 3 so you can use y = 3. That's if you use two separate equations. Does it need to be one equation?

OpenStudy (anonymous):

yeah thats what it askes for and what one would that be like slope- intercept form or point-slope, or standard form

OpenStudy (anonymous):

standard but if you need one equation that won't work

OpenStudy (anonymous):

so what do I do then?

OpenStudy (anonymous):

???

OpenStudy (anonymous):

or would I do it like y=3 and x=6? Taht is the 2 equations right?

OpenStudy (anonymous):

Since the line crosses x at 6 you can use the coordinate of (6,0) and since y crosses the at 3 you can use the point (0,3). Now with those two points find slope.

OpenStudy (anonymous):

What do you get for slope?

OpenStudy (anonymous):

idk i just asked this and idk the answer>> Write an equation in slope-intercept form, point-slope, or standard form for the line with the given information. Explain why you chose the form you used. Has an x-intercept of 6 and a y-intercept of 3

OpenStudy (anonymous):

slope subtract the y's 3-0 then subtract the x's: 0-6. You would have 3/-6 which reduces to -1/2. Now use the slope-intercept form. Y = -1/2x now find your y. Since is said it crosses the y at 3 your y is a positive 3. So your answer is y = -1/2x + 3

OpenStudy (anonymous):

Wow thanks!!!!!!!!!!!! One mroe??